# Hierarchical Scheduling for Multi-Vector Image Retrieval

### v1

@@ IV-A Redundancy Analysis with Retrieval Granularity
Specifically, we compute similarity scores with image segments in each granularity hierarchy iteratively. In each iteration (hierarchy), images with low aggregated similarity scores are pruned and will not be computed in the following granularities. The procedure is elaborated in L.2–L.6 of Alg. 1. The ratio of images pruned in each granularity hierarchy is specified with $t_{g}=\alpha^{g}T$, where $T$ denotes the initial reduction ratio and $\alpha$ denotes the decay rate. Empirically, $T$ can be more aggressive, as ground truth usually ranks near the top. While it is better to set a conservative $\alpha$, in that the score rank of ground truth may fluctuate as the granularity becomes finer.
Algorithm 1 Online Redundancy Reduction
1:procedure ProcessQuery($\{q_{k}\}$)
2:Let $\operatorname{Score}[N_{D},N_{q},N_{G}]$, $\operatorname{TopK}[N_{G}]$
3:for $g:=1$ to $N_{G}$ do
4:  for $k\in[1,N_{q}]$, $i\in[1,N_{D}t_{g-1}]$, $j\in[1,N_{g}]$ do
5:   $\operatorname{Score}_{i,k,g}\leftarrow\max(\operatorname{SIM}(q_{k},D_{i,j}^{g}),\operatorname{Score}_{i,k,g-1})$
[…]
7:  $\operatorname{TopK}[g]:=D[1:N_{K}]$
8:  if $\operatorname{K\tau}(\text{TopK}[g],\text{TopK}[g-1])\geq\tau$ then
9:   break
10:return $\operatorname{TopK}[g]$
IV-C Early Retrieval Convergence $\rightarrow$ Hierarchy Depth Opt.
Hierarchy distribution analysis reveals that a substantial portion of queries do not align with the finest granularity. For these queries, matching against the finest-grained image segments is unnecessary and introduces excessive redundancy. To address this issue, we design a hierarchy skipping mechanism that monitors the confidence level across granularities. By terminating the search early once a confidence threshold is reached, the mechanism adaptively avoids redundant fine-grained computations, incurring only minimal accuracy loss.
@@
The key challenge is to define the confidence indicator. According to Eq. 3, similarity scores remain unchanged until one of the sub-queries finds a better match. Therefore, it is natural to stop when the top-$k$ result list stabilizes. To quantify this convergence, we employ Kendall’s $\tau$ coefficient [15], denoted $\operatorname{K\tau}$, which measures the ordinal association between two ranked sequences. A preliminary study (Fig. 4(c)) confirms that $\operatorname{K\tau}$ consistently converges across datasets. In other words, when $\operatorname{K\tau}$ reaches a steady value, finer-grained segments no longer improve the result. The procedure is detailed in Alg. 1, Lines 7–-9, where $\textit{TopK}[g]$ denotes the set of $N_{K}$ images with the highest similarity scores up to granularity $g$.
IV-D Retrieval Granularity Sparisty $\rightarrow$ Hierarchy Sparisty Opt.
Beyond depth redundancy, cross-hierarchy analysis further indicates that certain intermediate granularities rarely align optimally with image objects. Thus they introduce additional redundancy, since it is impractical to make skipping choices online without prior knowledge of the dataset. Motivated by this insight, we propose an offline hierarchy pruning algorithm that eliminates redundant granularities based on dataset profiling. This reduces retrieval cost without sacrificing accuracy, since best-aligned objects of pruned granularities can typically be captured in adjacent ones.
@@
Alg. 2 (L.1–L.12) details the algorithm. We construct the hierarchy by evenly distributing levels over $[1,N]$ with interval $S_{G}$, where a smaller $S_{G}$ yields higher flexibility but more redundancy. The algorithm iteratively removes the hierarchy with the least accuracy loss. To ensure scheduling flexibility, consecutive removals are prohibited to prevent scenarios where only fine-grained hierarchies remain, which limits opportunities for online optimizations.
V Framework Integration & Automation
@@
Based on the redundancy theory introduced in Section IV, we implement the HiMIR framework with an automated configuration algorithm for optimizing parameters towards various trade-off objectives.
V-A Automated Configuration
@@
HiMIR can flexibly adapt to different dataset characteristics and deployment scenarios, whether accuracy- or performance-oriented, by exposing a set of tunable parameters: the initial ratio $T$ and decay rate $\alpha$ for redundant matching optimization, the hierarchy skipping threshold $\tau$, and the granularity initialization stride $S_{G}$, as detailed in Section IV.
@@
$\operatorname{Latency}=N_{q}\sum\nolimits_{g=1}^{N_{G}}N_{g}\times t_{g}\times N_{D},$ (4)
@@
Since similarity computation is lightweight, the latency of matching each sub-query with an image segment can be treated as a constant and is omitted here for simplicity. Constrained by latency, we first establish the granularity hierarchy, which requires more extensive exploration (L.15). Subsequently, other parameters are tuned for the highest accuracy under each latency constraint as (L.16–L.17). The latency-guided search enables efficient automated configuration. With proper initialization, the entire procedure completes within tens of minutes on datasets in our experiments.
Algorithm 2 Automated Configuration
1:procedure SetGran($S_{G},\tau,\alpha,T$)
2:repeat
3:  $N_{g}:=N_{g}+S_{G};\ \{N_{g}\}:=\{N_{g}\}\cup N_{g}$
4:until $\text{Eval}(\{N_{g}\},\tau,\alpha,T)$ Converge to $\mathcal{A}$
5:repeat
6:  for $g:=1$ to $|\{N_{g}\}|$ do
7:   $\{Acc\}:=\{Acc\}\cup\text{Eval}(\{N_{g}\}-N_{g},\tau,\alpha,T)$
8:  repeat
9:   $g:=\operatorname{argmin}(\{\text{Acc}\})$; $\{\text{Acc}\}:=\{Acc\}-g$
10:  until $g$ not near the last removed granularity
11:  $\{N_{g}\}:=\{N_{g}\}-g$
12:until $\text{Eval}(\{N_{g}\})\leq\mathcal{A}$
13:procedure Configure($R_{\tau},R_{S},R_{\alpha},R_{T},R_{L}$)
14:for $L,S,\tau,\alpha,T$ in $R_{L},R_{S},R_{\tau},R_{\alpha},R_{T}$ do
15:  Let $\{N_{g}\}:=\textsc{SetGran}(S,R_{\tau},R_{\alpha},R_{T})$
16:  if $L\geq\operatorname{Latency}(\tau,T,\alpha,\{N_{g}\})$ and $\text{Cfg[L].Acc}<\operatorname{Eval}(\tau,T,\alpha,\{N_{g}\})$ then
17:   $\text{Cfg}[L]:=\{\tau,T,\alpha,\{N_{g}\}\}$
V-B Implementation Detail
In this section, we reveal details of our implementation. We added about 1k lines of code based on the open-source code from  [4] built on PyTorch [16] and Faiss [12]. For image decomposition, given that the boundaries of segments generated by SLIC are irregular, we pad the remaining area with a black background to the bounding box of each segment to form valid patches. For query decomposition, we deployed a local large language model using vLLM [17], where the prompt template was adopted directly from [4]. For query processing, we optimize performance by vectorizing and parallelizing CPU-GPU operations with minimal memory transferring overhead.
VI Experiments
TABLE I: Main Results
Datasets Crepe MS Coco NoCaps Flickr Average
@@ VI Experiments
VI-A Experiment Setup
Evaluation Platform. As our framework is hardware-agnostic, we evaluate it on a standard platform equipped with an Intel Xeon 4410T CPU and one NVIDIA A100-PCIE-40G GPU.
Baselines. We compare our scheduling framework against the following two baseline methods:
@@ VI-A Experiment Setup
  • POQD [4]: a state-of-the-art multi-vector retrieval framework matching decomposed query with image segments, corresponding to the “1+N Mode” in Eq. 2.
Note that the proposed techniques are designed specifically for retrieval acceleration. Therefore, we do not include the LLM-based query splitting process in our performance comparison.
Datasets. Extending the setup of [4], we evaluate on 4 text-image datasets: CREPE [18], MsCoco [19], NoCaps [20], and Flickr [21], with 1K, 40K, 2K, and 2K images respectively, after data cleaning. Considering the scale of MSCOCO, a subset is randomly sampled for experiments. Note that we treat each image caption as a query and consider a caption relevant only to its corresponding image, following [18].
Models. Embedding and decomposition models are critical to retrieval accuracy. For text-image aligned embedding, we employ CLIP [6] to embed both queries and images. For query decomposition, we use Qwen3-8B [22] as the default large language model, in place of the model used in [4], owing to its improved decomposition performance.
Metrics. To evaluate efficiency, we measure the query throughput (queries per second, QPS) of HiMIR and baselines. For retrieval accuracy, we report the Normalized Discounted Cumulative Gain (NDCG [23]) at top-1 and top-10, denoted as NDCG@1 and NDCG@10.
Figure 6: Trade-off Analysis: Redundant Matching Opt.
VI-B Experiment Results
The main experimental results are summarized in Tab. I. We set four configurations to demonstrate the effectiveness of each proposed techniques: (1) only “1+M+N” algorithm without redundancy reduction. (2) hierarchy pruning (O3), (3) redundant matching optimization(O2), and (4) hierarchy depth optimization (O1). The parameter setting is obtained through the automated configuration framework. Note that we mainly compare our framework with MVR systems (POQD baseline), and thus the average speedup and accuracy improvement of the Vanilla baseline is omitted.
@@ VI-B Experiment Results
We further analyze the trade-offs of individual techniques. Fig. 6 illustrates the cross-hierarchy search. The curve shows throughput while the bar shows accuracy. As expected, filtering more images at each granularity (i.e., reducing $T$ and $\alpha$) significantly improves throughput. However, computation does not always translate to accuracy. For example, setting $\alpha$=0.8 yields the highest accuracy on Crepe, surpassing $\alpha$=1.0. It is common in CREPE that some images share common fine-grained objects but differ in composition. Thus hierarchical filtering helps eliminate such interference. This phenomenon is less evident in MS COCO, since Crepe has higher visual complexity. Fig. 7 investigates hierarchy skipping. The trend is consistent across datasets: larger granularity budgets $\tau$ improve accuracy by considering more granularities for finer query–image alignment, at the expense of throughput. This again highlights the intuitive performance–accuracy trade-off.
In summary, although more computation generally favors accuracy, the trade-off between accuracy and performance is non-trivial. This necessitates our automatic configuration framework. As shown in Fig. 8, HiMIR consistently pushes the Pareto frontier. POQD is represented by the green dot at the lower-left corner, while Vanilla dense retrieval lies outside the figure due to its extremely high throughput but poor accuracy.
Figure 7: Trade-off Analysis: Hierarchy Depth Opt.
Figure 8: Pareto Frontier Analysis
VI-C Overhead Analysis
The scheduling overhead of HiMIR consists of two parts. The first is the runtime early-exit metric computation, which compares the ranking of top candidates across consecutive iterations. This step has constant time complexity. The second is the additional sorting per iteration. Since the number of granularities is small (fewer than 10), the overall overhead is negligible, less than 0.1 ms per query.
VII Conclusion
This paper introduced HiMIR, a hierarchical decomposition framework for multi-vector image retrieval. By extending conventional “1+N Mode” retrieval into “1+M+N Mode,” HiMIR adapts to diverse image granularities while systematically exploiting redundancy and sparsity to improve efficiency. Through this co-design of algorithm, computation, and automation, HiMIR achieves significant gains in both accuracy and runtime cost, making fine-grained multimodal retrieval practical for real-world deployment. HiMIR opens up opportunities for broader integration with multimodal LLM systems and for further optimization of specific applications.

### v2

@@ 4.2. Low-similarity Tail Pruning
Specifically, we compute similarity scores for image segments in each granularity hierarchy and update the aggregated scores for images iteratively. In each iteration, images with low aggregated similarity scores are pruned and excluded from computation in subsequent hierarchies, as shown in Fig. 5(a). The procedure is elaborated in L.2–L.7 of Alg. 1. The ratio of images pruned in each granularity hierarchy is specified with $t_{g}=\alpha^{g}T$, where $T$ denotes the initial reduction ratio and $\alpha$ denotes the decay rate. Empirically, $T$ can be more aggressive, as ground truth usually ranks near the top. While it is better to set a conservative $\alpha$, in that the score rank of ground truth may fluctuate as the granularity becomes finer.
4.3. Hierarchy Depth Optimization
While Sec. 4.2 optimizes redundant fine-grained computation from the image side, we observe that similar opportunities also exist from the query side. Depth distribution in Fig. 4(b) shows that a substantial portion of queries align with coarse-grained information, indicating that they do not require traversing the full hierarchy. For these queries, matching against the finest-grained segments introduces redundant computation. To address this, we develop an early-exit mechanism that monitors confidence as granularity becomes finer and exits once it exceeds a predefined threshold.
@@ 4.3. Hierarchy Depth Optimization
The key challenge is to define the confidence indicator. A natural choice is to stop once the top-$k$ list stabilizes. To quantify this convergence, we employ Kendall’s $\tau$ coefficient (kendall1938new), denoted $\operatorname{K\tau}$, which measures the ordinal association between two ranked sequences. As illustrated in Fig. 5(b), we compute $\operatorname{K\tau}$ for the top-$k$ rankings between consecutive iterations as a confidence measure and halt as soon as it reaches a certain threshold, based on the intuition that a stabilized $\operatorname{K\tau}$ indicates ranking convergence. The procedure is detailed in Alg. 1 (L.8–L.10), where $\textit{TopK}[g]$ denotes the set of $N_{K}$ images with the highest similarity scores up to granularity $g$.
4.4. Hollow Hierarchy Elimination
Hollow hierarchy analysis further indicates that some intermediate hierarchies are unnecessary as they substantially overlap with their neighbors. However, making skip choices online is impractical without any prior knowledge, and pre-defining the hierarchical granularity set lacks adaptability. Therefore, we treat the granularity set as a tunable parameter and propose an offline hierarchy elimination algorithm to remove redundant hierarchies. Specifically, we run the algorithm on a validation query split while the image database remains unchanged, allowing us to optimize the hierarchy based on performance observed on this split.
@@ 4.4. Hollow Hierarchy Elimination
Alg. 2 (L.1–L.12) details the algorithm. We construct the hierarchy by initializing granularities spaced at intervals of $S_{G}$, where a smaller $S_{G}$ yields higher flexibility but more redundancy. The algorithm iteratively removes the granularity with the least accuracy loss before reaching a threshold. To ensure scheduling flexibility, consecutive removals are prohibited to prevent scenarios where only fine-grained hierarchies remain, limiting opportunities for online optimizations. Fig. 5(c) shows an example of this approach, where the granularity of 9 is considered non-essential and removed from the hierarchy, thereby reducing computational overhead.
5. Framework Integration & Automation
Algorithm 2 Automated Configuration
1:procedure SetGran($S_{G},\tau,T,\alpha$)
2:   $G:=0$; $\{N_{g}\}:=\emptyset$
3:   repeat
4:   $G:=G+S_{G};\ \{N_{g}\}:=\{N_{g}\}\cup G$ $\triangleright$ Init hierarchy with stride $S_{G}$
5:   until $\operatorname{EvalAcc}(\{N_{g}\},\tau,T,\alpha)$ converge to $\mathcal{A}$
6:   repeat
7:   for $i:=1$ to $|\{N_{g}\}|$ do
8:    $Acc_{i}:=\operatorname{EvalAcc}(\{N_{g}\}-N_{i},\tau,T,\alpha)$
9:   $g_{\text{remove}}:=\operatorname{argmax}(\{Acc_{i}\mid N_{i}$ not near the last removed one})\})
10:   $\{N_{g}\}:=\{N_{g}\}-N_{g_{\text{remove}}}$
11:   until $\operatorname{EvalAcc}(\{N_{g}\},\tau,T,\alpha)\leq\mathcal{A}-\epsilon$
12:   return $\{N_{g}\}$
13:procedure Configure($R_{L},R_{S},R_{\tau},R_{T},R_{\alpha}$)
14:   for $L,S,\tau,T,\alpha$ in $R_{L},R_{S},R_{\tau},R_{T},R_{\alpha}$ do
15:   Let $\{N_{g}\}:=\textsc{SetGran}(S,\tau,T,\alpha)$
16:   if $L\geq\operatorname{Latency}(\{N_{g}\},\tau,T,\alpha)$ and $\operatorname{EvalAcc}(\text{Cfg}[L])<\operatorname{EvalAcc}(\{N_{g}\},\tau,T,\alpha)$ then
17:    $\text{Cfg}[L]:=\{\{N_{g}\},\tau,T,\alpha\}$
5.1. Automated Configuration
@@
Based on the redundancy theory introduced in Section 4, we implement MIRAGE with an automated configuration algorithm for optimizing parameters for various trade-off objectives.
@@
MIRAGE can flexibly adapt to different dataset characteristics and deployment scenarios, whether accuracy or performance oriented, by exposing a set of tunable parameters: the initial ratio $T$ and decay rate $\alpha$ for low-similarity image pruning, the hierarchy early-exit threshold $\tau$, and the granularity initialization stride $S_{G}$, as detailed in Section 4.
@@
(4) $\operatorname{Latency}=N_{q}\sum\nolimits_{g=1}^{N_{G}}N_{g}\times t_{g}\times N_{D},$
@@
Since similarity computation is lightweight, the latency of matching each sub-query with an image segment can be treated as a constant and is omitted for simplicity. Constrained by latency, we first establish the granularity hierarchy, which requires more extensive exploration (L.15). Subsequently, other parameters are tuned for the highest accuracy under each latency constraint as (L.16–L.17). The latency-guided search enables efficient automated configuration.
It is worth noting that, in most retrieval scenarios (dubey2021retrievalsurvey), the database remains relatively static and can therefore be pre-analyzed offline to improve retrieval efficiency. We adopt the same assumption and use a separate validation query split solely for parameter tuning. With proper initialization, the entire preprocessing procedure finishes within tens of minutes on the datasets used in our experiments, which is acceptable for practical deployment.
5.2. Implementation Detail
We added about 1.1k lines of Python code based on the open-source code from  (liu2025poqd) built on PyTorch (pytorch) and Faiss (douze2024faiss), implementing our proposed scheduling framework and automated configuration algorithm, where the latency is profiled with CUDA Event API. For image decomposition, given that the boundaries of segments generated by SLIC are irregular, we pad the remaining area with a black background to the bounding box of each segment to form valid patches. For query decomposition, we deploy a local large language model using vLLM (kwon2023efficient), and the prompt template is adopted directly from (liu2025poqd) for fair comparison.
6. Experiments
Table 1. Main Results
Datasets CREPE MSCOCO NoCaps Flickr Improvement(Avg.)
@@ 6. Experiments
6.1. Experiment Setup
Evaluation Platform. We evaluate our framework on an edge server-level platform equipped with an Intel Xeon 4410T CPU and one NVIDIA A100-PCIE-40G GPU.
Baselines. We compare our framework against two baselines:
@@ 6.1. Experiment Setup
  • POQD (liu2025poqd): a state-of-the-art multi-vector retrieval framework matching decomposed query with image segments, corresponding to the “1+N Mode” in Eq. 2.
Note that the proposed techniques are designed specifically for retrieval acceleration. Therefore, we do not include the LLM-based query decomposing process in our performance comparison.
Datasets. Extending the setup of (liu2025poqd), we evaluate our method and baselines on 4 text-image retrieval datasets: CREPE (ma2023crepe), MSCOCO (lin2014mscoco), NoCaps (agrawal2019nocaps), and Flickr (young2014image), with 1K, 40K, 2K, and 2K images, respectively, after data cleaning. Considering the scale of MSCOCO, a subset is randomly sampled for experiments. Note that we treat each image caption as a query and consider a caption relevant only to its corresponding image, following (ma2023crepe). A subset of the captions is then selected as the test query set.
Models. Embedding and decomposition models are critical to retrieval accuracy. For text-image aligned embedding, we employ CLIP (radford2021learning) to embed both queries and images. For query decomposition, we use Qwen3-8B (yang2025qwen3) in place of the model used in (liu2025poqd), owing to its improved decomposition performance.
Metrics. To evaluate efficiency, we measure the query throughput (queries per second, QPS) of MIRAGE and baselines. For retrieval accuracy, we report the Normalized Discounted Cumulative Gain (NDCG (wang2013ndcg)) at top-10, denoted as NDCG@10, following (liu2025poqd).
6.2. End-to-End Results
The main experimental results are summarized in Tab. 1. We set four configurations to demonstrate the effectiveness of each proposed technique: (1) only “1+M+N” algorithm without redundancy reduction. (2) hollow hierarchy elimination (O3), (3) low-similarity tail pruning (O2), and (4) hierarchy depth optimization (O1). The parameter setting is obtained through the automated configuration algorithm. Note that we mainly compare our framework with MVR systems (POQD baseline), and thus, the average speedup and accuracy improvement of the Vanilla baseline are omitted.
Figure 6. Trade-off Analysis: Low-Similarity Tail Pruning
A AA
@@ 6.2. End-to-End Results
6.3. Analysis
In this section, we further present analysis on the trade-offs and effectiveness of individual techniques.
Fig. 6 analyzes low-similarity tail pruning. The curve shows throughput while the bar shows accuracy. As expected, filtering more images at each hierarchy (i.e., reducing $T$ and $\alpha$) boosts throughput. However, computation does not always translate to accuracy. For example, setting $\alpha$=0.8 yields the highest accuracy on CREPE, surpassing $\alpha$=1.0. It is common in CREPE that some images share common fine-grained objects but differ in composition. Thus, filtering tail images helps eliminate such interference. This phenomenon is less evident in MSCOCO, since CREPE has higher visual complexity. Fig. 7 investigates hierarchy depth optimization. The trend is consistent across datasets: a larger early-exit threshold $\tau$ improves accuracy by considering more granularities for finer query–image alignment, at the expense of throughput. This again highlights the intuitive performance–accuracy trade-off.
Although more computation generally favors accuracy, the trade-off between accuracy and performance is non-trivial, thus necessitating automatic configuration. As shown in Fig. 8(a), MIRAGE consistently pushes the Pareto frontier. The bottom-left green dot represents POQD, while Vanilla lies outside the figure due to its extremely high throughput but poor accuracy. The blue dots represent the samples generated in our search space, and among these marked with orange color are candidates reaching the optimal trade-off.
Figure 7. Trade-off Analysis: Hierarchy Depth Opt.
A AA
Figure 8. Component Analysis
A AA
To validate the effectiveness of Kendall’s $\tau$ coefficient, we present the evolution of $\operatorname{K\tau}$ over iterations in Fig. 8(b). The results show that $\operatorname{K\tau}$ consistently converges across datasets, confirming its reliability as an indicator for early exit.
The scheduling overhead of MIRAGE consists of two parts. The first is the runtime early-exit metric computation, which compares the ranking of top candidates across consecutive iterations. This step has constant time complexity. The second is the additional sorting per iteration. Since the number of granularities is small, the overall overhead is negligible (<0.1 ms) per query.
7. Conclusion
This paper introduced MIRAGE, a hierarchical decomposition framework for multi-vector image retrieval. By extending conventional “1+N Mode” retrieval into “1+M+N Mode,” MIRAGE adapts to diverse image granularities while systematically exploiting computational redundancy to improve efficiency. Through this co-design of algorithm, computation, and automation, MIRAGE achieves significant gains in both accuracy and runtime cost, making fine-grained multimodal retrieval practical for real-world deployment. MIRAGE opens up opportunities for broader integration with multimodal LLM systems and for further optimization of specific applications.